PI: Venturi
Title: 0420-014, 2345-167

Session: June 1996 (session 2)
Frequency: 6 cm
Mode: E
Obs time (UT): 169 EU 02:00(17/06)-13:30(17/06)
Stations: JbNtOnWbToSmUrSh
Processor: Bonn
EU Mk3 Tapes/St.: 2


To give feedback on this experiment, click here

Experiment Feedback

Onsala (John Conway at Mon Jun 17 13:20:37 EDT 1996) Success
No known problems

Noto (Corrado Trigilio at Tue Jun 18 09:24:11 EDT 1996) Success
No problems reported

Westerbork (Tony Foley at Thu Jun 20 05:53:17 EDT 1996) Minor failures
Day number wrong! Formatter had day number 168 in stead of 169 throughout the observation. Day number did not increment at 0UT.

Jodrell (David Henstock at Fri Jun 21 09:08:57 EDT 1996) Minor failures
Problems mounting first tape and getting the equipment working properly (overheating?) meant that experiment did not start recording until 2.56 UT (i.e. first 56 minutes lost). First tape with good data is NSF00321 - originally tape HS001586 had been mounted buit this was discarded before the start @02:56 UT,and it is belived there is no data on that tape. Around 13:10 UT the telescope was off-source, and some data will have been lost.

Urumqi (Lee Suqin at Tue Jul 2 01:50:24 EDT 1996) Minor failures
It was cloudy; clock offset = 0.5 /-0.5us. Because no power, two hours data lost(02:00-03:51). Tsys=313K; effective factor =49.5%; cal temp =67K. No phase cal for 6cm receiver.

Simeiz (Alexandre Volvatch at Tue Jul 2 06:12:16 EDT 1996) Minor failures
3 passes (07:52-08:20) were omited because source was behind the hill. 13:17 pass was omited because we had problem with radiotelescope.

Shanghai (Xiao Yu Tan at Mon Jul 15 12:47:39 EDT 1996) Minor failures
Clock offset = -0.6 micro sec. Obs okay, only LCP. Weather was cloudy.

HartRAO (Jonathan Quick at Wed Jul 17 08:07:49 EDT 1996) Success
Tapes only arrived 25 June 1996. No other problems.

Calibration (Fredrik T Rantakyro at Thu Oct 3 11:53:57 EDT 1996) Minor failures
An ANTAB file for this experiment can be found on VLBEER in the JUN96 directory.